OpenStudy (nuccioreggie):

A scale drawing of a living room is shown below. The scale is 1 : 40. A rectangle is shown. The length of the rectangle is labeled 6 inches. The width of the rectangle is labeled 4 inches. Show your work to determine the area of the room in square feet.

OpenStudy (nuccioreggie):

Help please

OpenStudy (nuccioreggie):

This is wat i got for the answer the area of the rectangle is 4x6 and its 1/40th the size of the room 4 × 6 =24 24 =1/ 40x ( 24 )( 40)= x So then we multiply 24 × 40 = 960 I think its 80ft

OpenStudy (nuccioreggie):

@Michele_Laino

OpenStudy (horsegirl27):

You are correct. The area is 24, then multiplied by 40. Good job! Have any questions?